# Quantum Bit Threads and the Entropohedron

### v1

@@ 5.1 Definitions & basic properties
$F_{S}$ is a non-empty compact convex polyhedron in $\mathbf{R}^{{X}}$, and is invariant under $f\to-f$.
@@
The codimension of $F_{S}$ in $\mathbf{R}^{X}$ equals the number of disjoint non-empty subsets $A\subseteq X$ such that $S(A)=0$.
@@
Given entropy functions $\{S_{i}\}$ for $X$ and non-negative numbers $\{\alpha_{i}\}$,
@@
is an entropy function for $X$, and
$F_{S}=\sum_{i}\alpha_{i}F_{S_{i}}\,.$ (183)
For a product state, for which the entropy function is the direct sum, the entropohedron is the Cartesian product.
@@
Let $S_{1},S_{2}$ be entropy functions respectively on disjoint sets ${X}_{1},{X}_{2}$. If $S$ is the entropy function on ${X}={X}_{1}\cup{X}_{2}$ defined as follows,
$S(A_{1}\cup A_{2}):=S_{1}(A_{1})+S_{2}(A_{2})\qquad(A_{1}\in 2^{{X}_{1}},A_{2}\in 2^{{X}_{2}})\,$, (184)
then
@@
Figure 11: Entropohedra for various states on three parties: (a) maximally mixed. (b) Maximally classically correlated. (c) Maximally entangled (e.g. GHZ or W state). (d) Marginal of four-party perfect tensor. (Figures are scaled to have equal single-party entropies.)
Examples of entropohedra for various quantum states on two and three parties are shown in Figs. 10, 11.
@@ 5.2 Saturation & positivity
The boundary of $F_{S}$ is the locus where one or more of the inequalities (181) is saturated. One can therefore ask under what circumstances some inequalities can be saturated in a single EDF. One may also want to know whether an EDF can be made everywhere non-negative, or non-positive, on a given set of parties while saturating an inequality. The following theorem addresses these questions, showing that nesting is the key property for saturating multiple inequalities. Specifically, one can saturate the inequality on any nested set of subsets, while keeping the EDF non-negative on the smallest subset. In fact, one can do better, saturating the inequality with one sign on one nested set of subsets and with the other sign on another nested set of subsets disjoint from the first set.
@@
Let $S$ be an entropy function, and let $A_{1}\subset\cdots\subset A_{m}$ and $B_{1}\subset\cdots B_{n}$ be elements of $2^{X}$ such that $A_{m}\cap B_{n}=\emptyset$. Then there exists an EDF $f$ such that:
@@
If $A_{1},A_{2}$ are nested, then this is always true, but otherwise it need not be. This is why the $A_{i}$s must be nested. A similar argument shows that the $B_{j}$s must also be nested, and disjoint from the $A_{i}$s.
@@
The fact that $f$ may be made non-negative within any given subset $A_{1}$ while saturating on that subset makes it an example of an entanglement contour function Vidal:2014aal , which to some extent characterizes the distribution of entanglement between $A_{1}$ and its complement. However, our viewpoint here is that, more than any one EDF, it is the set of all EDFs — the entropohedron — that captures the pattern of entanglement in the given state.
@@
for all $A\in 2^{X}\setminus\{\emptyset\}$. (If $S(A)=0$ for some $A\neq\emptyset$, then the dual polyhedron extends to infinity in the corresponding direction in the dual space to $\mathbf{R}^{X}$.) Because every region $A$ can be saturated, i.e. there exist EDFs $f$ such that
$\sum_{i\in A}f(x)=\pm S(A)\,$, (189)
@@ 5.3 Information quantities
where $A,B$ are disjoint sets and $AB:=A\cup B$. Let $L_{1}$ be the locus in $F_{S}$ where $A$ and $AB$ are both saturated, which is generically a codimension-2 edge on the boundary of $F_{S}$. On $L_{1}$,
@@
Let $L_{2}$ be the locus where $B$ and $AB$ are saturated, also generically a codimension-2 edge of $F_{S}$. On $L_{2}$,
$\sum_{x\in A}f(x)=S(AB)-S(B)\,,\qquad\sum_{x\in B}f(x)=S(A)\,.$ (192)
$L_{1}$ and $L_{2}$ both adjoin the codimension-1 face of $F_{S}$ where $AB$ is saturated; in fact, they are opposite edges of that face. They occupy parallel planes in $\mathbf{R}^{X}$, defined by constancy of the same linear functionals, namely $\sum_{x\in A}f(x)$ and $\sum_{x\in B}f(x)$. The relative displacement between those planes is
@@
where $\vec{v}(A)$ is the vector in $\mathbf{R}^{X}$ with $x$-component 1 if $x\in A$ and 0 if $x\notin A$. Note that, whereas the scalar quantity $I(A:B)$ is symmetric in its arguments, the vector $\vec{I}(A:B)$ is antisymmetric.
@@
I(A:C|B):=S(AB)+S(BC)S(B)S(ABC),I(A:C|B):=S(AB)+S(BC)-S(B)-S(ABC)\,, (194)
where $A,B,C$ are disjoint. We define $L_{1}$ as the locus in $F_{S}$ where $B,AB,ABC$ are all saturated, and $L_{2}$ as the locus where $B,BC,ABC$ are all saturated. On $L_{1}$ we have
$\sum_{x\in A}f(x)=S(AB)-S(B)\,,\qquad\sum_{x\in B}f(x)=S(B)\,,\qquad\sum_{x\in C}f(x)=S(ABC)-S(AB)\,$, (195)
@@
$\sum_{x\in A}f(x)=S(ABC)-S(BC)\,,\qquad\sum_{x\in B}f(x)=S(B)\,,\qquad\sum_{x\in C}f(x)=S(BC)-S(B)\,.$ (196)
These are generically codimension-3 edges of $F_{S}$. They are opposite edges of the codimension-2 face of $F_{S}$ where $B,ABC$ are saturated. They occupy parallel planes in $\mathbf{R}^{X}$, which are displaced relative to each other by the vector
$\vec{I}(A:C|B)=I(A:C|B)\left(\vec{v}(A)-\vec{v}(C)\right).$ (197)
@@ 5.4.1 Adding & removing parties, purification
If ${X}^{\prime}\subset{X}$, then the restriction of an entropy function $S$ for $X$ to subsets of ${X}^{\prime}$ is an entropy function $S^{\prime}$ for $X^{\prime}$. Furthermore, the restriction to ${X}^{\prime}$ of any EDF $f$ for $S$ is an EDF for $S^{\prime}$.
@@
An application of this corollary is to purification. The entropy function $S$ on X{X} is said to purify the entropy function $S^{\prime}$ on ${X}^{\prime}\subset{X}$ if S(X)=0S({X})=0 and, for all $A\subseteq X^{\prime}$, $S(A)=S^{\prime}(A)$. Given ${X}^{\prime}$ and $S^{\prime}$, such a purification can be constructed by adding a single element $O$ to ${X}^{\prime}$, so ${X}={X}^{\prime}\cup\{O\}$, and defining $S$ as follows:
@@
Since S(X)=0S({X})=0, for any EDF on $S$ we must have $\sum_{x\in{X}}f(x)=0$. Given an EDF $f^{\prime}$ for $S^{\prime}$, there therefore exists a unique extension $f$ for $S$, setting $f(O)=-\sum_{x\in{X}^{\prime}}f(x)$. Geometrically, the entropohedron $F_{S^{\prime}}\subset\mathbf{R}^{{X}^{\prime}}$ gets lifted onto the hypersurface $\sum_{x}f(x)=0$ in $\mathbf{R}^{{X}}$ to give $F_{S}$. Removing from X{X} any element of ${X}^{\prime}$ to obtain a set ${X}^{\prime\prime}$ then projects $F_{S}$ down to $F_{S^{\prime\prime}}\subset\mathbf{R}^{{X}^{\prime\prime}}$. All of these maps are one-to-one (since they are linear and dimension-preserving), so no information is lost. This is illustrated in Fig. 12.
@@ 5.4.3 Partial minimization
It’s clear that $S^{\prime}(\emptyset)=0$. So we just need to show that $S^{\prime}$ obeys (179), (180). Letting $\hat{A}_{\rm min}$, $\hat{B}_{\rm min}$ be minimizers for S(AA^)S($\cup\hat{A}$), S(BB^)S($\cup\hat{B}$) respectively, we have
@@
For the graph min cut function (see subsection 5.5), $S^{\prime}$ is the min cut function obtained by converting the boundary vertices in $X\setminus X^{\prime}$ into internal (i.e. non-boundary) vertices. In the quantum setting, we do not know what operation or channel (if any) acting on the state on the parties $X$ with entropies $S$ gives the state on the parties $X^{\prime}$ with entropies $S^{\prime}$.
@@
5.5 Relation to flows on graphs
Entropy functions can be derived not just from quantum states but also from graphs. Specifically, given a weighted undirected graph, with a subset X{X} of the vertex set labelled as “boundary” vertices, the min cut function satisfies (178)–(180). (It defines a “pure state” in the sense that S(X)=0S({X})=0.) For a given $A\in 2^{X}$, by the max flow-min cut theorem, $S(A)$ equals the maximal flux through the graph from $A$ to ${X}\setminus A$. As we will show in theorem 5.15 below, there is a direct relationship between the set of flows and the entropohedron: A point in $f\in\mathbf{R}^{X}$ is in $F_{S}$ if and only if it equals the boundary flux of some flow. This is the graph version of theorem 2.2 from section 2.
Not every entropy function can be realized as the min cut function on a graph. The latter obey additional inequalities, such as $I_{3}(A:B:C)\geq 0$ (where the negative tripartite information $I_{3}$ was defined in (198)). The set of entropy functions obtainable from graphs, for a given number of parties, is called the holographic entropy cone Bao:2015bfa . Its structure for a general number of parties remains unknown (see Grimaldi:2025jad for an overview and references). It would be interesting if the entropohedron might have some special properties for such entropy functions, and if it could be used to shed some light on the holographic entropy cone.
Before formally stating theorem 5.15, we remind the reader of the relevant definitions and basic facts about graphs. A weighted undirected graph is defined by a set $I$ of internal vertices, a set X{X} of external (or boundary) vertices, and a set $E$ of edges, together with a positive function $w$ on $E$. Each edge $e$ is a pair of distinct vertices. A cut $r$ is a subset of $I\cup{X}$, its edge set $\eth r$ is the set of edges with one vertex inside and the other outside $C$, and its weight $|\eth r|$ is the total weight of its edge set:
@@ 5.5 Relation to flows on graphs
Although the graph is undirected, in order to define flows we need to assign an arbitrary fiducial orientation to them; we then define a sign $s(i,e)=\pm 1$ for $i\in e$, with $s(i,e)=+1$ if the orientation of $e$ points toward the vertex $i$ and s(i,e)=1s(i,e)=-1 if it points away. A flow on the graph is a function $v:E\to\mathbf{R}$ obeying
@@
We are now ready to state and prove theorem 5.15:
@@
Given a weighted undirected graph with boundary vertex set X{X}, let $S$ be the corresponding min cut function on $2^{X}$. A point $f\in\mathbf{R}^{X}$ is in $F_{S}$ if and only if $f=\Phi_{v}$ for some flow $v$.
@@
It follows directly from the definitions that, for any flow $v$, $\Phi_{v}\in F_{v}$. The more non-trivial statement is the converse: for any $f\in F_{S}$, there exists a flow $v$ such that $\Phi_{v}=f$. To prove this, we set up the following convex program:
@@
Such a convex program with objective 0 is a feasibility test: If there exists a feasible flow, then the optimal value is 0; if not, it is $\infty$. Dualizing the program, we will see which one is correct.
@@
By the assumption that $f\in F_{S}$, the first term in the integrand is bounded above by $)S(A(\hat{\lambda}$)). The second term is bounded above by $)-S(A(\hat{\lambda}$)) by the definition of $S$. So the total is non-positive for all $\hat{\lambda}$. Therefore, the maximum is 0, achieved by $\lambda=0$. (As a check, note that, if $f$ violates the EDF condition for some $A\in 2^{X}$, then by making $\lambda$ constant on the minimal cut for $A$ and 0 elsewhere, the integral can be made arbitrarily large and positive. Therefore, the maximum is $\infty$, so as expected the primal program does not admit a feasible point.) ∎

### v2

@@ 5.2 Basic properties
Given an entropy function $S$ for $X$, the entropohedron $F_{S}$ is a non-empty compact convex polyhedron in $\mathbf{R}^{{X}}$, and is invariant under $f\to-f$.
@@
Given an entropy function $S$ for $X$, the codimension of the entropohedron $F_{S}$ in $\mathbf{R}^{X}$ equals the maximum number of disjoint non-empty sets $A\subseteq X$ such that $S(A)=0$.
@@
Given a set of pairs $\{(\alpha_{i},S_{i})\}$, where $\alpha_{i}$ is a non-negative number and $S_{i}$ is an entropy function for $X$, the function $S:2^{X}\to\mathbf{R}$ defined by
@@
is an entropy function, and its entropohedron is
$F_{S}=\sum_{i}\alpha_{i}F_{S_{i}}:=\left\{\sum_{i}\alpha_{i}f_{i}:f_{1}\in F_{S_{1}},f_{2}\in F_{S_{2}},\ldots\right\}.$ (184)
Under direct sums of entropy functions (e.g. for tensor products of quantum states), the entropohedron is the Cartesian product.
@@
Given entropy functions $S_{1},S_{2}$ on disjoint sets ${X}_{1},{X}_{2}$ respectively, the function $S:2^{X_{1}\cup X_{2}}\to\mathbf{R}$ defined by
$S(A_{1}\cup A_{2}):=S_{1}(A_{1})+S_{2}(A_{2})\qquad(A_{1}\subseteq{{X}_{1}},A_{2}\subseteq{{X}_{2}})$ (185)
is an entropy function, and its entropohedron is
@@ 5.3 Saturation & positivity
The boundary of $F_{S}$ is the locus where one or more of the defining inequalities (182) is saturated. One can therefore ask under what circumstances some inequalities can be saturated in a single EDF. One may also want to know whether an EDF can be made everywhere non-negative, or non-positive, on a given set of parties while saturating an inequality. The following theorem addresses these questions, showing that nesting is the key property for saturating multiple inequalities. Specifically, one can saturate the inequality on any nested set of subsets, while keeping the EDF non-negative on the smallest subset. In fact, one can do better, saturating the inequality with one sign on one nested set of subsets and with the other sign on another nested set of subsets disjoint from the first set.
@@
Let $S$ be an entropy function for $X$, and let $A_{1}\subset\cdots\subset A_{m}$ and $B_{1}\subset\cdots\subset B_{n}$ be subsets of $X$ such that $A_{m}\cap B_{n}=\emptyset$. Then there exists an EDF $f$ such that:
@@
If $A_{1},A_{2}$ are nested, then this is always true, but otherwise it need not be. This is why the theorem requires the $A_{i}$s to be nested. A similar argument shows that the $B_{j}$s must also be nested, and disjoint from the $A_{i}$s.
@@
Specialized to the case $m=1$, $n=0$, theorem 5.5 gives us a function $f$ that, within $A_{1}$, is non-negative and sums to $S(A_{1})$. This makes it an example of an entanglement contour Vidal and Chen (2014). The purpose of the entanglement contour is to characterize the distribution of entanglement between $A_{1}$ and its complement. However, our viewpoint here is that, more than any one EDF, it is the set of all EDFs — the entropohedron — that best captures the distribution of entanglement in the given state.
@@
for all non-empty $A\subseteq X$. (If $S(A)=0$ for some non-empty $A$, then the dual polyhedron extends to infinity in the corresponding direction in the dual space to $\mathbf{R}^{X}$.) Because every region $A$ can be saturated, i.e. there exist EDFs $f$ such that
$\sum_{x\in A}f(x)=\pm S(A)\,$, (190)
@@ 5.4 Information quantities
Figure 12: Geometric manifestation of information quantities on the entropohedron: (a) The entropohedron for a correlated pair of bits (as shown in Fig. 10(b)); with $A=\{1\}$, $B=\{2\}$, we show the loci $L_{A}$, $L_{B}$ and their relative displacement, the vector $\vec{I}(A:B)$ (red), which quantifies the mutual information (see (194)). (b) Two views of the entropohedron for the marginal of a four-party perfect tensor (as shown in Fig. 21(d)); with $A=\{1\}$, $B=\{2\}$, $C=\{3\}$, we show the loci $L_{A}$, $L_{B}$ and their relative displacement,, the vector $\vec{I}(A:B|C)$ (blue), which quantifies the conditional mutual information (see (198)).
@@
where $A,B$ are disjoint sets and $AB:=A\cup B$. Let $L_{A}$ be the locus in $F_{S}$ where $A$ and $AB$ are both saturated, which is generically a codimension-2 edge on the boundary of $F_{S}$. On $L_{A}$,
@@
Let $L_{B}$ be the locus where $B$ and $AB$ are saturated, also generically a codimension-2 edge of $F_{S}$. On $L_{B}$,
$\sum_{x\in A}f(x)=S(AB)-S(B)\,,\qquad\sum_{x\in B}f(x)=S(B)\,.$ (193)
$L_{A}$ and $L_{B}$ both adjoin the codimension-1 face of $F_{S}$ where $AB$ is saturated; in fact, they are opposite edges of that face. They sit in parallel planes in $\mathbf{R}^{X}$, defined by constancy of the same linear functionals, namely $\sum_{x\in A}f(x)$ and $\sum_{x\in B}f(x)$. The relative displacement between those planes is
@@
where $\vec{v}(A)$ is the vector in $\mathbf{R}^{X}$ with $x$-component 1 if $x\in A$ and 0 if $x\notin A$. See Fig. 12(a) for an example. Note that, whereas the scalar quantity $I(A:B)$ is symmetric in its arguments, the vector $\vec{I}(A:B)$ is antisymmetric.
@@
I(A:B|C):=S(AC)+S(BC)S(C)S(ABC),I(A:B|C):=S(AC)+S(BC)-S(C)-S(ABC)\,, (195)
where $A,B,C$ are disjoint. We define $L_{1}$ as the locus in $F_{S}$ where $C,AC,ABC$ are all saturated, and $L_{2}$ as the locus where $B,BC,ABC$ are all saturated. On $L_{1}$ we have
$\sum_{x\in A}f(x)=S(AC)-S(C)\,,\qquad\sum_{x\in B}f(x)=S(ABC)-S(AC)\,,\qquad\sum_{x\in C}f(x)=S(C)\,$, (196)
@@
$\sum_{x\in A}f(x)=S(ABC)-S(BC)\,,\qquad\sum_{x\in B}f(x)=S(BC)-S(C)\,,\qquad\sum_{x\in C}f(x)=S(C)\,.$ (197)
These are generically codimension-3 edges of $F_{S}$. They are opposite edges of the codimension-2 face of $F_{S}$ where $C,ABC$ are saturated. They sit in parallel planes in $\mathbf{R}^{X}$, which are displaced relative to each other by the vector
$\vec{I}(A:B|C)=I(A:B|C)\left(\vec{v}(A)-\vec{v}(B)\right).$ (198)
See Fig. 12(b) for an example.
@@ 5.5.1 Adding & removing parties, purification
If $S$ is an entropy function for $X$ and ${X}^{\prime}\subset{X}$, then the restriction of $S$ to subsets $A\subseteq X^{\prime}$ is an entropy function $S^{\prime}$ for $X^{\prime}$. Furthermore, the restriction to ${X}^{\prime}$ of any EDF $f$ for $S$ is an EDF for $S^{\prime}$.
@@
An application of this corollary is to purification. The entropy function $S$ on X{X} is said to purify the entropy function $S^{\prime}$ on ${X}^{\prime}\subset{X}$ if S(X)=0S({X})=0 and, for all $A\subseteq X^{\prime}$, $S(A)=S^{\prime}(A)$. Given ${X}^{\prime}$ and $S^{\prime}$, a minimal purification can be constructed by adding a single element $O$ to ${X}^{\prime}$, so ${X}={X}^{\prime}\cup\{O\}$, and defining $S$ as follows:
@@
Since S(X)=0S({X})=0, for any EDF on $S$ we must have $\sum_{x\in{X}}f(x)=0$. Given an EDF $f^{\prime}$ for $S^{\prime}$, there therefore exists a unique extension $f$ for $S$, setting $f(O)=-\sum_{x\in{X}^{\prime}}f(x)$. Geometrically, the entropohedron $F_{S^{\prime}}\subset\mathbf{R}^{{X}^{\prime}}$ gets lifted onto the “purity” hyperplane
$\left\{f\in\mathbf{R}^{X}:\sum_{x\in X}f(x)=0\right\}$ (202)
to give $F_{S}$. Removing from X{X} any element of ${X}^{\prime}$ to obtain a set ${X}^{\prime\prime}$ then projects $F_{S}$ down to $F_{S^{\prime\prime}}\subset\mathbf{R}^{{X}^{\prime\prime}}$. All of these maps are bijective (since they are linear and dimension-preserving), so no information is lost in the process. This is illustrated in Fig. 13.
@@ 5.5.3 Partial minimization
It’s clear that $S^{\prime}(\emptyset)=0$. To show that $S^{\prime}$ obeys (180), let $\hat{A}_{\rm min}$, $\hat{B}_{\rm min}$ be minimizers for S(AA^)S($\cup\hat{A}$), S(BB^)S($\cup\hat{B}$) respectively. We have
@@
For the graph min cut function (see subsection 5.6), $S^{\prime}$ is the min cut function obtained by converting the boundary vertices in $X\setminus X^{\prime}$ into internal (i.e. non-boundary) vertices. This shows that partial minimization maps the holographic entropy cone (defined in that subsection) to itself. In the quantum setting, we do not know what kind of operation, if any, acting on the state on the parties $X$ with entropies $S$ gives the state on the parties $X^{\prime}$ with entropies $S^{\prime}$. [^22] An interesting question in this context is whether partial minimization preserves the constrained inequalities valid for quantum states Linden and Winter (2005); Cadney et al. (2012).
@@
5.6 Relation to network flows
Entropy functions can be derived not just from quantum states but also from networks. Specifically, given a weighted undirected graph, with a subset X{X} of the vertex set labelled as “external” vertices, the min cut function on $X$ satisfies (179)–(181). (This function defines a “pure state” in the sense that S(X)=0S({X})=0.) For a given set $A\subseteq X$, by the max flow-min cut theorem, $S(A)$ equals the maximal flux through the graph from $A$ to ${X}\setminus A$. As we will show in theorem 5.15 below, there is a direct relationship between the set of flows and the entropohedron: A point $f\in\mathbf{R}^{X}$ is in $F_{S}$ if and only if it equals the boundary flux of some flow. This is the graph version of theorem 2.2 from section 2.
Not every entropy function can be realized as the min cut function on a network. Network min cut functions obey additional inequalities beyond (180), (181), such as $I_{3}(A:B:C)\geq 0$ (where the negative tripartite information $I_{3}$ was defined in (199)). The set of entropy functions obtainable from networks, for a given number of parties, is called the holographic entropy cone Bao et al. (2015). Its structure for a general number of parties remains unknown (see Grimaldi et al. (2025) for an overview and references). It would be interesting if the entropohedron might have some special properties for such entropy functions, and if it could be used to shed some light on the holographic entropy cone.
Before formally stating theorem 5.15, we remind the reader of the relevant definitions and facts about graphs. We’ll use the term network to refer to a weighted undirected graph with some vertices labelled as “external” and the rest as “internal”. Specifically, a network $(X,I,E,w)$ consists of a set X{X} of external vertices, a set $I$ of internal vertices, a set $E$ of edges, and a positive function $w$ on $E$. Each edge $e\in E$ is an unordered pair of distinct vertices. A cut $r$ is a subset of $I\cup{X}$, its edge set $\eth r$ is the set of edges with one vertex inside and the other outside $C$, and its weight $|\eth r|$ is the total weight of its edge set:
@@ 5.6 Relation to network flows
Although the graph is undirected, in order to define flows we need to assign an arbitrary fiducial orientation to the edges; we then define a sign $s(i,e)=\pm 1$ for $i\in e$, with $s(i,e)=+1$ if the orientation of $e$ points toward the vertex $i$ and s(i,e)=1s(i,e)=-1 if it points away. A flow on the graph is a function $v:E\to\mathbf{R}$ obeying
@@
We are now ready to state and prove the relation between flows and the min cut entropohedron.
@@
Given a network $(X,I,E,w)$, let $S$ be the corresponding min cut function on $2^{X}$. A point $f\in\mathbf{R}^{X}$ is in $F_{S}$ if and only if $f=\Phi_{v}$ for some flow $v$ on the network.
@@
It follows directly from the definitions that, for any flow $v$, $\Phi_{v}\in F_{S}$. The more non-trivial statement is the converse: for any $f\in F_{S}$, there exists a flow $v$ such that $\Phi_{v}=f$. To prove this, we set up the following convex program:
@@
Such a convex program with objective 0 is a feasibility test: If there exists a feasible flow, then the optimal value is 0; if not, it is $\infty$. Dualizing the program, we will see which one is correct. Since the program can be written as a linear program, strong duality automatically holds.
@@
By the assumption that $f\in F_{S}$, the first term in the integrand is bounded above by $)S(A(\hat{\lambda}$)). The second term is bounded above by $)-S(A(\hat{\lambda}$)) by the definition of $S$. So the total is non-positive for all $\hat{\lambda}$. Therefore, the maximum is 0, achieved by $\lambda=0$. (As a check, note that, if $f$ violates the EDF condition for some $A\subseteq X$, then by making $\lambda$ constant on the minimal cut for $A$ and 0 elsewhere, the integral can be made arbitrarily large and positive. Therefore, the maximum is $\infty$, so as expected the primal program does not admit a feasible point.) ∎